# Teaching Large Language Models an Unseen Language on the Fly

### v1

@@ 5.3 Additional Experiments on Kalamang
It is extremely hard to identify a language completely unseen by current LLMs and collect enough resources for it. Besides ZhuangBench, the only suitable evaluation dataset is from a contemporary work Tanzer et al. (2024), which presents a dataset called MTOB. It consists of translation tasks between English (eng) and Kalamang (kgv), another low-resource language unseen by current LLMs.
@@
We report preliminary results on MTOB in Table 2. DiPMT++ outperforms the baseline in the original paper of MTOB across most settings. This further proves that DiPMT++ is a language-agnostic framework and can adapt to different low-resource languages without extra efforts. See more details in Appendix E.
@@
GPT-3.5-Turbo 2.9 35.5 11.4 37.0
Table 2: Results of the original baseline from Tanzer et al. (2024) and DiPMT++ on the test set of MTOB. The best scores are made bold, with the second underlined.
@@ 6 Discussion
Here we delve into two important questions when adapting LLMs to an unseen language. One is how to improve the coverage of lexical explanations given limited or incomplete resources. We carefully analyze the three strategies used in DiPMT++ to alleviate the out-of-dictionary problem. The other is how to teach LLMs syntactical rules solely through prompting. We go beyond DiPMT++ and investigate more potential strategies to help LLMs learn syntax implicitly or explicitly.
@@ 6.1 Improving Lexical Coverage
During the translation, one might not find all the words in the dictionary. For example, with the original dictionary in ZhuangBench, we can only find the meaning for 67% of the Zhuang words and 47% of the Chinese words in the test set of ZhuangBench. As described in Section 4.2, we adopt three strategies to improve the lexical coverage of the prompt. Here we analyze how they contribute to the performance of DiPMT++.
@@
Different strategies can address different types of out-of-dictionary problems. We qualitatively analyze the types of words recalled by each strategy on ZhuangBench. Fuzzy matching greatly helps address the compound words in Chinese and Zhuang. For example, the out-of-dictionary Chinese word 日常生活 (daily life) is decomposed into 日常 (daily) and 生活 (life) by forward maximal matching. Similarly, the Zhuang word mehlwg (mother and child) is decomposed into meh (mother) and lwg (child). By bilingual lexicon induction on our corpus, we can mine common words overlooked by the dictionaries, i.e., soujgih $\Leftrightarrow$ 手机 (mobile phone), or newly-invented words, i.e., lienhcanjnieb $\Leftrightarrow$ 产业链 (industry chain). Since these strategies are language-agnostic, they can also be applied to other languages and alleviate the out-of-dictionary problems caused by certain morphological phenomena.
@@ 6.2 Learning Syntax
For extremely low-resource languages like Zhuang, there is neither a large-scale monolingual corpus available for language modeling nor well-defined formal grammar such as context-free grammar (CFG) ready to use. Given the strong reasoning ability of LLMs, one might wonder to what extent LLMs can learn the syntax of an unseen language, either implicitly or explicitly, through prompting. We conduct a preliminary study of different strategies for teaching LLMs syntax. From our positive or negative results, we reveal what strategies work and what might not currently.
@@ 6.2.1 Implicit Learning
Learn from DiPMT++ Exemplars. In DiPMT++, we retrieve exemplars from the corpus with BM25, a language-agnostic retrieval algorithm, hoping LLMs can infer shallow syntax from them. Besides BM25, we try two other strategies. One is randomly sampling exemplars. The other is retrieval based on part-of-speech (POS) sequence [^6] As there are no POS taggers available for Zhuang, the POS sequence of a Zhuang sentence is approximated by concatenating the POS tag of each word’s Chinese translation., assuming that sentences with similar POS sequences (measured by Levenshtein distance) may share similar syntactical structures.
As shown in Table 4, the exemplars retrieved by BM25 greatly improve the translation quality over the random sampling, as they contain richer syntactic information related to the current testing instance. POS-based retrieval might not provide much assistance to the model’s syntactic learning, as the abstraction from natural language sentences to POS sequences is overly simplified and the POS sequences are sometimes noisy.
@@
Learn from Monolingual Texts. Inspired by learning language modeling from a large-scale corpus, we wonder whether LLMs can quickly familiarize themselves with the syntax of an unseen language through a small piece of monolingual text in the prompt, when we have only a few thousand tokens of texts for a low-resource language. Therefore, we add monolingual Zhuang texts to the DiPMT++ prompt, hoping that it can help LLMs generate more coherent Zhuang sentences.
We test this strategy with GPT-3.5-Turbo [^7] We initially test this strategy using Qwen-14B-chat. Qwen-14B-chat fails to produce meaningful outputs when presented with an additional 1K tokens of Zhuang texts in the prompt. Incorporating large portions of texts in an unseen language might overwhelm this medium-sized LLM.. Table 5 shows a 0.9 BLEU gain in zh2za translation after adding 1K Zhuang tokens. Notably, the easy subset sees a substantial 2.8-point increase. As the length of the monolingual text increases, we do not observe a continuous increase in performance. We find that when provided with 5K monolingual text, the model becomes overconfident in this language and fabricates more words that are not present in the prompt. We refrain from adding longer monolingual texts than 5K for the high expense.
@@
Table 5: BLEU scores of adding different lengths of monolingual texts to the prompt of DiPMT++, using GPT-3.5-Turbo on the zh2za task of ZhuangBench.
@@ 6.2.2 Explicit Learning
We choose the order of modifiers and modified elements as our research focus. Different from Chinese and English, Zhuang usually places modifiers after modified elements. Our pilot study shows presenting this syntactical rule directly within the prompt proves to be largely ineffective, as it might be overly abstract for LLMs to grasp. When instructed to adopt this rule to analyze this specific syntactical phenomenon in a chain-of-thought style, LLMs demonstrate a better capacity to manage it, successfully addressing 70% of the studied cases. See Appendix F for details.
@@ 7 Assisting Human Translation
Here we discuss the potential of applying DiPMT++ to realistic scenarios. Through a pilot user study, we show that we can use LLMs to assist humans in understanding an extremely low-resource language, even if both the LLMs and the humans have no prior knowledge of this language.
@@ 7.1 Study Design
We carry out a user study to investigate to what extent an LLM can help humans translate a completely unseen language.
Settings. We compare three settings: (1) LLM Only: An LLM is prompted with DiPMT++ and outputs a translation. (2) Human Only: We ask humans to use the given linguistic resources (i.e., a dictionary and a corpus of parallel sentences) to perform the translation. (3) Human + LLM: We provide humans with an LLM for assistance, in addition to linguistic resources. Humans can refer to the initial translation results from the LLM.
Data & Model. From the easy subset of ZhuangBench, we sample 20 instances for zh2za and 40 instances for za2zh. We use Qwen-14B-chat for experiments. This open-source model performs decently on ZhuangBench and has a medium size of parameters, with affordable computational cost.
Participants. We recruit 6 graduate students majoring in NLP, who are speakers of Chinese but have no prior knowledge of Zhuang. Each instance is translated by at least 4 participants. For a given instance, some participants are asked to translate it with LLM assistance while others without it.
See details of the study in Appendix G.
@@ 7.2 Results and Analysis
Time normal-$\downarrow$ Score normal-$\uparrow$ Time normal-$\downarrow$ Score normal-$\uparrow$
@@
Quality and Efficiency. Table 6 shows the study results. Surprisingly, with the provided linguistic resources, the participants can properly translate simple sentences written in an unseen language into their native language. Despite investing much time, the average BLEU scores of their translations are 10+ points higher than those of LLM.
Providing initial translation outputs from the LLM yields enhancements in human translation quality. For zh2za translation, the LLM helps increase the human performance by 2.5 BLEU while the improvement is 2.9 BLEU for za2zh translation.
Furthermore, the LLM greatly boosts the efficiency of zh2za translation. The participants save 17% of their time on average, as they can leverage the LLM’s outputs rather than crafting translations from scratch. For za2zh translation, we observe no obvious difference in the efficiency between the two settings. It is probably because in the Human Only setting, the participants, who are native Chinese speakers, excel in identifying plausible Chinese words from the given prompt and structuring them into coherent sentences. This process requires less time compared to meticulously verifying the LLM-generated output.
Human Actions. During za2zh translation, the participants perform an average of 2.1 dictionary searches and 1.3 corpus searches for each instance. Conversely, for the harder zh2za translation, we observe a higher frequency of searches: an average of 3.5 dictionary searches and 5.4 corpus searches.
We note that many participants exhibit a pattern of switching between searches conducted in two languages, aiming to find $n$-gram evidence to support the translation of specific words or phrases. See examples in Appendix G. This strategy inherently aligns with the retrieval-augmented generation (RAG) framework, which involves alternating between retrieval and generation stages. Such observations offer insights into innovative solutions for on-the-fly language learning. We will release the human action sequences for future research.
Broader Applications. Besides aiding humans in translation, the LLMs enhanced with the DiPMT++ framework have broader applications for low-resource languages. These include education for underrepresented languages, preservation of endangered languages, and research into historical or extinct languages. We anticipate that by these techniques, researchers can contribute to the preservation of linguistic diversity worldwide.
8 Conclusion
In this paper, we investigate whether LLMs can learn a completely new language on the fly. Our experiments on the Zhuang language show that LLMs can rapidly grasp an unseen language through proper ICL. However, challenges persist in analyzing more intricate morphological phenomena and achieving a more comprehensive and precise understanding of syntax. We hope that ZhuangBench and DiPMT++ can encourage more research on efficient methods for underrepresented languages.
@@ 8 Conclusion
Scale of Evaluation. Due to the limited language resources, the evaluation scale is relatively small. Our evaluation is based on only 200 Zhuang-Chinese testing instances and 50 Kalamang-English testing instances. We plan to expand the size of the testing set.
Typology of Studied Languages. Despite belonging to different language families (Krai-Dai and Sino-Tibetan, respectively), Zhuang and Chinese share many similarities. Like Chinese, Zhuang exhibits few inflectional morphologies and has many loanwords from Chinese. The high degree of similarity between Zhuang and Chinese may lead to overly optimistic conclusions. More research on languages with larger differences, such as that we conduct on Kalamang and English, would provide a more comprehensive understanding.
Scope of Studied Methods. Our exploration of explicitly learning syntactic information is limited to analyzing a specific syntactical phenomenon with CoT reasoning. Other potential methods, such as using external grammar books might may be more suitable for more powerful LLMs. We did not do it due to budget constraints. Otherwise, more comprehensive conclusions could have been obtained.
@@ Appendix A The Zhuang Language
Zhuang is a group of Kra–Dai languages spoken by the Zhuang people of Southern China in the province of Guangxi and adjacent parts of Yunnan and Guangdong. It is used by more than 16 million people. The current official writing system for Zhuang is the Latin script. Zhuang is considered as an isolating language with few inflectional morphology. In our work, we focus on Standard Zhuang, the official standardized form of the Zhuang language.
Zhuang has been largely overlooked in current NLP research, evidenced by the absence of an open-source corpus dedicated to the language. The lack of available training data means that popular open-source multilingual models like mBERT Devlin et al. (2019), BLOOM Workshop et al. (2022), and NLLB Costa-jussà et al. (2022) do not include any support for Zhuang. Even competitive commercial models like GPT-3.5 and GPT-4 exhibit near-zero proficiency when they comes to Zhuang. This lack of support underscores the challenges faced in developing NLP solutions for low-resource languages like Zhuang.
@@ Appendix B Dataset Collection and Statistics
Data Checking.
@@
Statistic of Parallel Corpus.
@@
Statistic of Translation Test Set.
@@ Appendix C Implementation Details
Prompt Construction.
In Table 10 we provide an example of the prompt for Zhuang-to-Chinese translation used in DiPMT++. We use 3-shot exemplars, as our preliminary experiments show that using three exemplars can achieve decent performance while having affordable computational cost. For each word in the source sentence, we provide two possible meanings from the dictionary.
Dictionary Expansion.

### v2

@@ 5.3 Additional Experiments on Other Languages
It is extremely hard to identify a language completely unseen by current LLMs and collect enough resources for it. Besides ZhuangBench, the only suitable evaluation dataset is MTOB, from a contemporary work Tanzer et al. (2024). It consists of translation tasks between English (eng) and Kalamang (kgv), another low-resource language unseen by current LLMs.
@@
We report preliminary results on MTOB in Table 2. DiPMT++ outperforms the baseline in the original paper of MTOB across most settings. This further proves that DiPMT++ is a language-agnostic framework and can adapt to different low-resource languages without extra effort. See details in Appendix E.
@@
GPT-3.5 2.9 35.5 11.4 37.0
Table 2: Results of the original baseline from Tanzer et al. (2024) and DiPMT++ on the test set of MTOB. The original baseline is the W + S setting in Tanzer et al. (2024). The best scores are made bold, with the second underlined.
Seen Languages
We also evaluate DiPMT++ on 7 low-resource languages that might have been seen during pre-training. Unlike unseen languages such as Zhuang and Kalamang, LLMs can translate these languages with a non-zero BLEU score by zero-shot prompting. DiPMT++ can still improve the translation quality for extremely low-resource languages, whose BLEU scores are below 10 originally. See details in Appendix F.
@@ 6 Discussion
Here we delve into two important questions when adapting LLMs to an unseen language. One is how to improve the coverage of lexical explanations given limited or incomplete resources. We analyze the three strategies used in DiPMT++ to alleviate the out-of-dictionary problem. The other is how to teach LLMs syntactic rules solely through prompting. We go beyond DiPMT++ and investigate more potential strategies to help LLMs learn syntax implicitly or explicitly.
@@ 6.1 Improving Lexical Coverage
During the translation, one might not find all the words in the dictionary. For example, with the original dictionary in ZhuangBench, we can only find the entries for 67% of the Zhuang words and 47% of the Chinese words in the test set of ZhuangBench. As described in Section 4.2, we adopt three strategies to improve the lexical coverage of the prompt. Here we analyze how they contribute to the performance of DiPMT++.
@@
[…]
@@ 6.2 Learning Syntax
For extremely low-resource languages like Zhuang, there is neither a large-scale monolingual corpus available for language modeling nor well-defined formal grammar such as context-free grammar (CFG) ready to use. Given the strong reasoning ability of LLMs, one might wonder to what extent LLMs can learn the syntax of an unseen language, either implicitly or explicitly, through prompting. We conduct a preliminary study of different strategies for teaching LLMs syntax and reveal what strategies work and what might not currently.
@@ 6.2.1 Implicit Learning
Learn from DiPMT++ Exemplars
In DiPMT++, we retrieve exemplars from the corpus with BM25, a language-agnostic retrieval algorithm, hoping LLMs can infer shallow syntax from them. Besides BM25, we try two other strategies. One is randomly sampling exemplars. The other is retrieval based on part-of-speech (POS) sequence [^6] As there are no POS taggers available for Zhuang, the POS sequence of a Zhuang sentence is approximated by concatenating the POS tag of each word’s Chinese translation., assuming that sentences with similar POS sequences (measured by Levenshtein distance) may share similar syntactical structures.
As shown in Table 4, the exemplars retrieved by BM25 greatly improve the translation quality over the random sampling, as they contain richer lexical and syntactic information related to the testing instance. POS-based retrieval might not provide much assistance to the model’s syntactic learning, as the abstraction from natural language sentences to POS sequences is overly simplified and the POS sequences are sometimes noisy.
@@
Learn from Monolingual Texts
Inspired by learning language modeling from a large-scale corpus, we wonder whether LLMs can quickly familiarize themselves with the syntax of an unseen language through a small piece of monolingual texts in the prompt when we have only a few thousand tokens of texts for a low-resource language. Therefore, we add monolingual Zhuang texts to the DiPMT++ prompt [^7] In this study, we did not design sophisticated strategies for the selection of monolingual texts. We directly extract monolingual texts of varying lengths from a news story in Zhuang, instead of retrieving texts for different testing instances. In the future, we will collect more monolingual corpora and explore clever strategies for selecting proper monolingual texts., hoping that it can help LLMs generate more coherent Zhuang sentences.
We test this strategy with GPT-3.5888We initially test this strategy using Qwen-14B-chat. This model fails to produce meaningful outputs when presented with an additional 1K tokens of Zhuang texts in the prompt. Incorporating large portions of texts in an unseen language might overwhelm this medium-sized LLM.. Table 5 shows a 0.9 BLEU gain in zh2za translation after adding 1K Zhuang tokens. Notably, the easy subset sees a substantial 2.8-point increase. As the length of the monolingual text increases, we do not observe a continuous increase in performance. We find that when provided with 5K monolingual text, the model becomes overconfident in this language and fabricates more words not present in the prompt. We refrain from adding longer monolingual texts than 5K for the high expense.
@@
Table 5: BLEU scores of adding different lengths of monolingual texts to the prompt of DiPMT++, using GPT-3.5 on the zh2za task of ZhuangBench.
@@ 6.2.2 Explicit Learning
We choose the order of modifiers and modified elements as our research focus. Different from Chinese and English, Zhuang usually places modifiers after modified elements. We select 10 zh2za testing samples where GPT-3.5 fails on the order of modifiers and modified elements, and try different strategies to incorporate this syntactic rule into the prompt.
First, we attempt to directly declare a rule of word order between modifiers and modified elements in the prompt, i.e., different from Chinese, Zhuang puts the modifiers after the modified elements. GPT-3.5 correctly addresses only 1 of the 10 testing examples.
Second, we use chain-of-thought (CoT) reasoning to encourage the model to analyze the word order in the given source sentence. The format of CoT is shown in Table 14. We notice that GPT-3.5 often fails to identify the modifiers and the modified elements from the given sentence, probably because of a lack of general linguistic knowledge. After we hint the modifiers and the modified elements of the given sentences in the prompt, GPT-3.5 can correctly analyze and address their order for 7 out of 10 testing examples.
Our pilot study demonstrates that understanding grammar rules is a complex procedure and still presents significant challenges to current LLMs. Although several works Tanzer et al. (2024); Gemini Team (2024) claim that LLMs are able to read grammar books with larger context lengths, it is still questionable whether they truly understand the grammar rules in the books.
@@ 7 Assisting Human Translation
Here we discuss the potential of applying DiPMT++ to realistic scenarios. Through a user study, we show that we can use LLMs to assist humans in understanding an extremely low-resource language, even if both the LLMs and the humans have no prior knowledge about this language.
@@ 7.1 Study Design
Our user study aims to investigate to what extent an LLM can help humans translate a completely unseen language.
Settings
We compare three settings: (1) LLM Only: An LLM is prompted with DiPMT++ and outputs a translation. (2) Human Only: We ask humans to use the given linguistic resources (i.e., a dictionary and a corpus of parallel sentences) to perform the translation. (3) Human + LLM: We provide humans with an LLM for assistance, in addition to the linguistic resources. Humans can refer to the initial translation results from the LLM.
Data & Model
From the easy subset of ZhuangBench, we sample 20 instances for zh2za and 40 for za2zh. We use Qwen-14B-chat for experiments. This open-source model performs decently on ZhuangBench and has a medium size of parameters with affordable computational cost.
Participants
We recruit 6 graduate students majoring in NLP, who are native speakers of Chinese but have no prior knowledge of Zhuang. The participants are not given training materials regarding the Zhuang language before carrying out the task. We only train them to use the interface through a few demonstrations. Each testing instance is translated by at least 4 participants. For a given testing instance, some participants are asked to translate it with LLM assistance while others are without it.
See details of the study in Appendix H.
@@ 7.2 Results and Analysis
Time $\downarrow$ Score $\uparrow$ Time $\downarrow$ Score $\uparrow$
@@
Quality and Efficiency
Table 6 shows the study results. Surprisingly, with the provided linguistic resources, the participants can properly translate simple sentences written in an unseen language into their native language. Despite costing much time, the average BLEU scores of their translations are 10+ points higher than those of the LLM.
Providing initial translation output from the LLM yields improvement in human translation quality. For zh2za translation, the LLM helps increase the human performance by 2.5 BLEU while the improvement is 2.9 BLEU for za2zh translation.
Furthermore, the LLM greatly boosts the efficiency of zh2za translation. The participants save 17% of their time on average, as they can leverage the LLM’s output rather than crafting translations from scratch. For za2zh translation, we observe no obvious difference in terms of efficiency between the two settings. It is probably because in the Human Only setting, the participants, who are native Chinese speakers, excel in identifying plausible Chinese words from the given prompt and structuring them into coherent sentences. This process requires less time than meticulously verifying the LLM-generated output.
Human Actions
During the za2zh translation, the participants perform an average of 2.1 dictionary searches and 1.3 corpus searches for each testing instance. Conversely, for the harder zh2za translation, we observe more frequent searches: an average of 3.5 dictionary searches and 5.4 corpus searches.
We note that many participants exhibit a pattern of switching between searches conducted in two languages, aiming to find $n$-gram evidence to support the translation of specific words or phrases. See examples in Appendix H. This strategy aligns with the retrieval-augmented generation (RAG) framework, which involves alternating between retrieval and generation stages. Such observations offer insights into innovative solutions for on-the-fly language learning.
Broader Applications
Besides aiding humans in translation, the LLMs enhanced with the DiPMT++ framework have broader applications for low-resource languages. These include education for underrepresented languages, preservation of endangered languages, and research into historical or extinct languages. We anticipate that by these techniques, researchers can better contribute to the linguistic diversity worldwide.
8 Conclusions
In this paper, we investigate whether LLMs can learn a completely new language on the fly. Our experiments on the Zhuang language show that LLMs can rapidly grasp an unseen language through proper ICL. However, challenges persist in analyzing more intricate morphological phenomena and achieving a more comprehensive and precise understanding of syntax. We hope that ZhuangBench and DiPMT++ can encourage more research efforts on efficient methods for underrepresented languages.
@@ 8 Conclusions
Scale of Evaluation
Due to the limited language resources, the evaluation scale is relatively small. Our evaluation is based on only 200 Zhuang-Chinese testing instances and 50 Kalamang-English testing instances. We plan to expand the size of the testing set.
Typology of Studied Languages
Despite belonging to different language families (Krai-Dai and Sino-Tibetan, respectively), Zhuang and Chinese share similarities. Like Chinese, Zhuang exhibits few inflectional morphologies and has many loanwords from Chinese. The similarities between Zhuang and Chinese may lead to overly optimistic conclusions. More research on languages with larger differences, such as that we conduct on Kalamang and English, would provide a more comprehensive understanding.
Scope of Studied Methods
Our exploration of explicitly learning syntactic information is limited to analyzing a specific syntactic phenomenon with CoT reasoning. Other potential methods, such as using external grammar books might may be suitable for more powerful LLMs. We did not do this due to budget constraints.
Acknowledgments
This work is supported in part by NSFC (62161160339) and Beijing Science and Technology Program (Z231100007423011). We thank the anonymous reviewers for their valuable suggestions. We thank Mingxu Tao, Zirui Wu, Zhibin Chen, and Quzhe Huang for their help in this work. For any correspondence, please contact Yansong Feng.
@@ Appendix A The Zhuang Language
Zhuang is a group of Kra–Dai languages spoken by the Zhuang people of Southern China in the province of Guangxi and adjacent parts of Yunnan and Guangdong. It is used by more than 16 million people. The current official writing system for Zhuang is the Latin script. Zhuang is considered an isolating language with little inflectional morphology. In this work, we focus on Standard Zhuang, the official standardized form of the Zhuang language.
Zhuang has been largely overlooked in current NLP research, evidenced by the absence of an open-source corpus dedicated to the language. The lack of available training data means that popular open-source multilingual models like mBERT Devlin et al. (2019), BLOOM Workshop et al. (2022), and NLLB Costa-jussà et al. (2022) do not include any support for Zhuang. Even competitive commercial models like GPT-3.5 and GPT-4 exhibit near-zero proficiency when it comes to Zhuang. This lack of support underscores the challenges faced in developing NLP solutions for low-resource languages like Zhuang.
@@ Appendix B Dataset Collection and Statistics
Collection Details
Here we explain how we collect parallel sentences from the official government reports. All the versions of the official government reports in other languages strictly correspond to the Chinese version, to ensure consistency in conveying information. The Chinese and Zhuang versions of the official reports have the same number of paragraphs, with each corresponding paragraph containing the same number of sentences. Therefore, we can achieve sentence-level mapping in a fully automated manner by segmenting the reports into sentences and then aligning them in sequence, without the need for manual annotation.
Data Checking
@@
Statistic of Parallel Corpus
@@
Statistic of Translation Test Set
@@ Appendix C Implementation Details
Prompt Construction
In Table 11 we provide an example of the prompt for Zhuang-to-Chinese translation used in DiPMT++. We use 3-shot exemplars, as our preliminary experiments show that using three exemplars can achieve decent performance while having affordable computational costs. For each word in the source sentence, we provide two possible meanings from the dictionary.
Dictionary Expansion
@@
Hyperparameters of Finetuning
For mT5, we adopt the seq2seq training script for machine translation by Transformers Wolf et al. (2020). The batch size is 4. The learning rate is 5e-5. We train the model for 3 epochs, evaluate the checkpoints every 500 steps and select the best-performing one.
For Llama-2-7B, we use the DeepSpeed framework Rasley et al. (2020). We use the instruction "Please translate the following Chinese sentence into Zhuang:" in the prompt. The batch size is 64. The learning rate is 2e-5. We train the model for 3 epochs, evaluate the checkpoints after each epoch and select the best-performing one.